OpenStudy (anonymous):

Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (mathstudent55):

Slope intercept method is rewrite the eq into this form: y = mx + b In other words, solve the equation for y 7x + 3y = 42 3y = -7x + 42 y = (-7/3)x + 14 Now the slope is m = -7/3, and the y intercept is 14 which means point (0,14)

OpenStudy (anonymous):

I have part 1 @mathstudent55 i need part 2

OpenStudy (mathstudent55):